If the consumer's budget constraint is given by 10F + 5S = 100 where F is food and S is shelter, how much food can he buy if he purchases 4 units of shelter?

A 10

B 8

C 20

D 9
Answer #1

Answer : The answer is option B.

Given,

10F + 5S = 100

And S = 4 units.

By putting the value of S in budget constraint we get,

10F + (5 * 4) = 100

=> 10F + 20 = 100

=> 10F = 100 - 20 = 80

=> F = 80 / 10

=> F = 8

So, here the consumer buy 8 units of food.

Therefore, option B is correct.